# Statistical-Computational Trade-offs for Greedy Recursive Partitioning Estimators

### v1

@@ Abstract
Models based on recursive partitioning such as decision trees and their ensembles are popular for high-dimensional regression as they can potentially avoid the curse of dimensionality. Because empirical risk minimization (ERM) is computationally infeasible, these models are typically trained using greedy algorithms. Although effective in many cases, these algorithms have been empirically observed to get stuck at local optima. We explore this phenomenon in the context of learning sparse regression functions over $d$ binary features, showing that when the true regression function $f^{*}$ does not satisfy the so-called Merged Staircase Property (MSP), greedy training requires $\exp(\Omega(d))$ to achieve low estimation error. Conversely, when $f^{*}$ does satisfy MSP, greedy training can attain small estimation error with only $O(\log d)$ samples. This performance mirrors that of two-layer neural networks trained with stochastic gradient descent (SGD) in the mean-field regime, thereby establishing a head-to-head comparison between SGD-trained neural networks and greedy recursive partitioning estimators. Furthermore, ERM-trained recursive partitioning estimators achieve low estimation error with $O(\log d)$ samples irrespective of whether $f^{*}$ satisfies MSP, thereby demonstrating a statistical-computational trade-off for greedy training. Our proofs are based on a novel interpretation of greedy recursive partitioning using stochastic process theory and a coupling technique that may be of independent interest.
@@ 1 Introduction
Decision tree models are piecewise constant supervised learning models obtained by recursive partitioning of the covariate space. Although classical, they remain among the most important supervised learning models because they are highly interpretable (Rudin et al., 2021; Murdoch et al., 2019) and yet are flexible enough to afford the potential for high prediction accuracy. This potential is maximized when decision trees are combined in ensembles via random forests (RFs) (Breiman, 2001) or gradient boosting (Friedman, 2001). These algorithms are widely recognized as having state-of-the-art performance on moderately-sized tabular datasets (Caruana et al., 2008; Fernández-Delgado et al., 2014; Olson et al., 2018), even outperforming state-of-the-art deep learning methods (Grinsztajn et al., 2022), despite the amount of attention lavished on the latter. Such datasets are common in many settings such as bioinformatics, healthcare, economics, and social sciences. Naturally, decision trees and their ensembles receive widespread use via their implementation in popular machine learning packages such as ranger (Wright et al., 2017), scikit-learn (Pedregosa et al., 2011), xgboost (Chen and Guestrin, 2016), and lgbm (Ke et al., 2017). Decision trees have also been adapted to a variety of tasks beyond regression and classification, including survival analysis (Ishwaran et al., 2008), heterogeneous treatment effect estimation (Athey and Imbens, 2016), time series analysis, and multi-task learning.
@@ 1.1 Consistency of ERM trees, Greedy Trees, and Non-Adaptive Trees
As a result, practitioners make use of greedy algorithms such as CART  (Breiman et al., 1984), ID3 (Quinlan, 1986), or C4.5 (Quinlan, 1993). Such algorithms grow a tree in a top-down manner, with the split for a given node chosen by optimizing a local node-specific objective. Once made, a split is never reversed. [^1] Decision trees are sometimes pruned as a one-off post-processing procedure after they are fully grown. These heuristics seem to work well in practice—both RFs and gradient boosting are built out of CART trees—but are notoriously difficult to analyze. Indeed, early efforts to prove consistency of greedy regression and classification trees under standard smoothness conditions on $f^{*}$ had to further assume that the bias of the tree converged to zero (Breiman et al., 1984; Chaudhuri et al., 1995). More recent works were able to replace this assumption with more reasonable alternatives on $f^{*}$ such as additivity  (Scornet et al., 2015; Klusowski, 2020; Klusowski and Tian, 2024), a submodularity condition (Syrgkanis and Zampetakis, 2020), or the sufficient impurity decrease (SID) condition (Chi et al., 2022; Mazumder and Wang, 2024).
@@ 1.2 High-Dimensional Consistency
We first argue that non-adaptive trees are not high-dimensional consistent [^2] See Appendix A for a formal definition of non-adaptive tree models and a rigorous proof of this statement.: Assuming relatively balanced splits, the average depth of a tree is $O(\log n)$. Hence, $\log n=o(d)$ implies that not every covariate can receive a split, or in other words, that the tree needs to be able to selectively split on relevant covariates, which contradicts the definition of non-adaptivity. In contrast, several works (Syrgkanis and Zampetakis, 2020; Chi et al., 2022; Mazumder and Wang, 2024; Klusowski and Tian, 2024) have showed that CART and RFs are high-dimensional consistent given various condition on $f^{*}$ (e.g., submodularity, SID, or additivity). This helps to explain the successful application of RFs to high-dimensional problems such as genomics in which $n\ll d$.
@@
Beyond some heuristic calculations, Syrgkanis and Zampetakis (2020) were not able to make this argument rigorous, nor did they attempt to generalize beyond this example.
@@
For any subset $S\subset[d]$ of the coordinate indices, let $\chi_{S}$ be the monomial function defined by $\chi_{S}(\mathbf{x})=\prod_{j\in S}x_{j}$. [^3] If $|S|=s$, such a function is sometimes called an $s$-sparse parity in the theoretical computer science literature. It is a classical result from Fourier analysis (Stein and Shakarchi, 2011) that every function $f^{*}\colon\{\pm 1\}^{d}\to\mathbb{R}$ can be uniquely written in the form
@@
for some nonnegative integer $r$, subsets $S_{1},S_{2},\ldots,S_{r}$ of $\left\{1,2,\ldots,d\right\}$ and nonzero coefficients $\alpha_{S_{j}},j=1,2,\ldots,d$. We say that $f^{*}$ satisfies the merged-staircase property (MSP) if the subsets $S_{1},S_{2},\ldots,S_{r}$ can be reordered such that for any $i\in\{1,2,\ldots,r\}$, we have
@@ 1.6 Robust Lower Bounds under “Soft” Bottlenecks
When $\mathcal{X}=\{\pm 1\}^{d}$, $\nu=\operatorname{Unif}\left(\{\pm 1\}^{d}\right)$, $f^{*}$ depends only on $s$ covariates, then the expected $L^{2}$ risk of the CART estimator $\hat{f}_{\operatorname{CART}}$ satisfies $E_{\mathcal{D}_{n}}\left\{R\left(\hat{f}(-;\mathcal{D}_{n}),f^{*}\right)\right\}=\Omega(1)$ whenever $n=\min\left\{O\left(1/\min_{f\notin\textsf{MSP}}R(f,f^{*})\right),\exp(O(d))\right\}$.
@@ 2.5 CART
where $\mathcal{I}(\mathcal{C};\mathcal{D}_{n})\coloneqq\frac{1}{N(\mathcal{C})}\sum_{i\in I(\mathcal{C})}\left(Y_{i}-\bar{Y}_{I(\mathcal{C})}\right)^{2}$ is called the impurity of the cell $\mathcal{C}$.
@@ 2.6 Random Forests
Slightly abusing notation, we denote the CART model fitted with a random seed $\theta$ as $\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n},\theta)$. A random forest model with $M$ trees is then
@@ 2.7 Greedy Tree Models
In order to isolate the aspect of CART that leads to its limitations with non-MSP functions, we make a more general definition. We say that $\mathcal{A}$ is a greedy tree algorithm if, similar to CART, it grows a tree greedily in a top-down fashion, but uses a possibly different criterion $\mathcal{O}(k;\mathcal{C},\mathcal{D}_{n})$ instead of impurity decrease (12) to determine its splits. We constrain $\mathcal{O}(k;\mathcal{C},\mathcal{D}_{n})$ to depend on $\mathcal{D}_{n}$ only via the values of the $k$-th covariate and the response for observations in $\mathcal{C}$, i.e.
@@ 7.1 Characterizing High-Dimensional Consistency for CART
Next, one can prove that $f^{*}$ is SMSP if and only if it satisfies the SID condition (7) for some value of $\lambda>0$ (see Proposition F.2). Previous work, specifically Chi et al. (2022) and Mazumder and Wang (2024), provide upper bounds for the estimation error of CART given the SID condition. These results and proofs are stated for continuous covariates, but can easily be adapted to a discrete covariate setting. We revisit this last statement in the next subsection. Assuming this for now, our discussion thus far yields the following characterization theorem.
@@
It is natural to ask whether our lower bounds can be extended to handle all $f^{*}$ that do not satisfy SMSP (or equivalently the SID condition). This turns out to be impossible—there exist functions that do not satisfy SMSP, but for which high-dimensional consistency holds. For instance, consider $f^{*}(\mathbf{x})=x_{1}+2x_{2}+2x_{1}x_{2}+x_{2}x_{3}$. Since $f^{*}|_{\left\{x_{1}=-1\right\}}=x_{2}x_{3}$, this does not satisfy SMSP, but on the other hand, CART would always first split on $x_{2}$ given enough samples, in which case this cancellation never occurs.
@@ 11 Discussion
Finally, while we have proved lower bounds for greedy axis-aligned regression trees, it is natural to ask whether lower bounds of a similar flavor hold for oblique regression trees as well as other more general classes of regression trees and ensembles under computational constraints. We hope our work will serve as a first step in clarifying the statistical-computational trade-offs inherent within the class of greedy algorithms fitting regressing trees and ensembles.
@@ Acknowledgments
The authors would like to thank Denny Wu, Wei-Yin Loh, Samory Kpotufe, Rahul Mazumder, and Yingying Fan for insightful discussions. Tan gratefully acknowledges support from NUS Start-up Grant A-8000448-00-00 and MOE AcRF Tier 1 Grant A-8002498-00-00. Klusowski gratefully acknowledges financial support from the National Science Foundation (NSF) through CAREER DMS-2239448. Balasubramanian gratefully acknowledges support from NSF through grants DMS-2053918 and DMS-2413426.
@@ Appendix I Constant Probability Lower Bounds
The lower bounds portion of Abbe et al. (2023)’s conjecture is suggested by existing Statistical Query (SQ) and Correlational Statistical Query (CSQ) lower bounds. SQ and CSQ are generic models of constrained computation (Kearns, 1998; Feldman et al., 2021). In the SQ model, the analyst has access to a sequence of queries, each of which returns an approximation of the expectation $\mathbb{E}\left\{\psi(\textbf{X},Y)\right\}$ for some real-valued function $\psi$ that may change from query to query. The CSQ model further constrains the queries to be of the form $\mathbb{E}\left\{\phi(\textbf{X})Y\right\}$ for some real-valued function $\phi$. Under the SQ model, it is known that $\Omega\left(d^{k}\right))$ queries are required to learn $k$-parities (Barak et al., 2022), while Abbe et al. (2023) proved a CSQ lower bound of the same order. Now, Abbe et al. (2021) showed that SGD is no more powerful than the SQ model if the stochastic gradients have limited precision, but it is unclear to us whether this holds for the SGD regime in 2LNNOSGD. Assuming the reduction holds, then the SQ and CSQ lower bounds translate to a lower sample complexity bound of $\tilde{\Theta}\left(d^{k}\right)$.
We also note an additional performance gap suggested by Abbe et al. (2023)’s conjecture. They in fact consider a generalization of sparse Boolean functions called multi-index functions. These are defined as $f^{*}(\mathbf{x})=g(\mathbf{M}\mathbf{x})$, where $\mathbf{M}$ is an unknown $d\times s$ matrix with orthogonal columns. In other words, $f^{*}$ depends only on an $s$-dimensional latent subspace. The fact that this subspace is not axis-aligned is irrelevant to SGD and neural networks, which are rotation-invariant. Glasgow (2024)’s upper bound for XOR was furthermore proved in this setting. [^7] Kou et al. (2024)’s result required axis alignment, which they exploited using sign SGD. On the other hand, having sparsity in the original covariate basis is crucial to the good performance of regression trees and ensembles in general. This is because every root-to-leaf path can split on at most $O(\log d)$ covariates, and hence cannot fit to any dense rotation of a latent subspace, no matter how small the latent dimension is (Tan et al., 2021). This drawback may be overcome using oblique (that is, non-axis-aligned) regression trees and ensembles (Cattaneo et al., 2024; O’Reilly, 2024), but these have much larger computational overhead. Finally, we note that for continuous covariates, the information exponent introduced by Arous et al. (2024) and its generalizations by Ren and Lee (2024) may be more effective in capturing learnability. Exploring similar measures of learnability for recursive partitioning trees and their ensembles presents an interesting direction for future research.

### v2

# Statistical-Computational Trade-offs for Recursive Adaptive Partitioning Estimators

@@ Abstract
Models based on recursive adaptive partitioning such as decision trees and their ensembles are popular for high-dimensional regression as they can potentially avoid the curse of dimensionality. Because empirical risk minimization (ERM) is computationally infeasible, these models are typically trained using greedy algorithms. Although effective in many cases, these algorithms have been empirically observed to get stuck at local optima. We explore this phenomenon in the context of learning sparse regression functions over $d$ binary features, showing that when the true regression function $f^{*}$ does not satisfy Abbe et al. (2022)’s Merged Staircase Property (MSP), greedy training requires $\exp(\Omega(d))$ to achieve low estimation error. Conversely, when $f^{*}$ does satisfy MSP, greedy training can attain small estimation error with only $O(\log d)$ samples. This dichotomy mirrors that of two-layer neural networks trained with stochastic gradient descent (SGD) in the mean-field regime, thereby establishing a head-to-head comparison between SGD-trained neural networks and greedy recursive partitioning estimators. Furthermore, ERM-trained recursive partitioning estimators achieve low estimation error with $O(\log d)$ samples irrespective of whether $f^{*}$ satisfies MSP, thereby demonstrating a statistical-computational trade-off for greedy training. Our proofs are based on a novel interpretation of greedy recursive partitioning using stochastic process theory and a coupling technique that may be of independent interest.
@@ 1 Introduction
Decision tree models are piecewise constant supervised learning models obtained by recursive adaptive partitioning of the covariate space. Although classical, they remain among the most important supervised learning models because they are highly interpretable (Rudin et al., 2021; Murdoch et al., 2019) and yet are flexible enough to afford the potential for high prediction accuracy. This potential is maximized when decision trees are combined in ensembles via random forests (RFs) (Breiman, 2001) or gradient boosting (Friedman, 2001). These algorithms are widely recognized as having state-of-the-art performance on moderately-sized tabular datasets (Caruana et al., 2008; Fernández-Delgado et al., 2014; Olson et al., 2018), even outperforming state-of-the-art deep learning methods (Grinsztajn et al., 2022), despite the amount of attention lavished on the latter. Such datasets are common in many settings such as bioinformatics, healthcare, economics, and social sciences. Naturally, decision trees and their ensembles receive widespread use via their implementation in popular machine learning packages such as ranger (Wright et al., 2017), scikit-learn (Pedregosa et al., 2011), xgboost (Chen and Guestrin, 2016), and lgbm (Ke et al., 2017). Decision trees have also been adapted to a variety of tasks beyond regression and classification, including survival analysis (Ishwaran et al., 2008), heterogeneous treatment effect estimation (Athey and Imbens, 2016), time series analysis, and multi-task learning.
@@ 1.1 Consistency of ERM trees, Greedy Trees, and Non-Adaptive Trees
As a result, practitioners make use of greedy algorithms such as CART  (Breiman et al., 1984), ID3 (Quinlan, 1986), or C4.5 (Quinlan, 1993). Such algorithms grow a tree in a top-down manner, with the split for a given node chosen by optimizing a local node-specific objective. Once made, a split is never reversed. [^1] Decision trees are sometimes pruned as a one-off post-processing procedure after they are fully grown. These heuristics seem to work well in practice—both RFs and gradient boosting are built out of CART trees—but are notoriously difficult to analyze. Indeed, early efforts to prove consistency of greedy regression and classification trees under standard smoothness conditions on $f^{*}$ had to further assume that the bias of the tree converged to zero (Breiman et al., 1984; Chaudhuri et al., 1995). More recent works were able to replace this assumption with more reasonable alternatives on $f^{*}$ such as additivity  (Scornet et al., 2015; Klusowski, 2020; Klusowski and Tian, 2024) or a submodularity condition (Syrgkanis and Zampetakis, 2020). Chi et al. (2022) further introduced and derived consistency under the sufficient impurity decrease (SID) condition, which we will define more formally later in this paper. Mazumder and Wang (2024) later improved upon the consistency rates of Chi et al. (2022) and showed that the SID condition is satisfied for larger classes of models.
@@ 1.2 High-Dimensional Consistency
We first argue that non-adaptive trees are not high-dimensional consistent [^2] See Appendix A for a formal definition of non-adaptive tree models and a rigorous proof of this statement.: Assuming relatively balanced splits, the average depth of a tree is $O(\log n)$. Hence, $\log n=o(d)$ implies that not every covariate can receive a split, or in other words, that the tree needs to be able to selectively split on relevant covariates, which contradicts the definition of non-adaptivity. In contrast, several works have showed that CART and RFs are high-dimensional consistent given various condition on $f^{*}$ (e.g., Syrgkanis and Zampetakis (2020) for binary covariates and assuming submodularity, Klusowski and Tian (2024) for arbitrary covariates and assuming additivity, Chi et al. (2022) for arbitrary covariates and assuming SID). This helps to explain the successful application of RFs to high-dimensional problems such as genomics in which $n\ll d$.
@@
Beyond some heuristic calculations, Syrgkanis and Zampetakis (2020) did not make this argument rigorous, nor did they generalize beyond the XOR example.
@@
For any subset $S\subset[d]$ of the coordinate indices, let $\chi_{S}$ be the monomial function defined by $\chi_{S}(\mathbf{x})=\prod_{j\in S}x_{j}$. [^3] If $|S|=s$, such a function is sometimes called an $s$-parity in the theoretical computer science literature. It is a classical result from Fourier analysis (Stein and Shakarchi, 2011) that every function $f^{*}\colon\{\pm 1\}^{d}\to\mathbb{R}$ can be uniquely written in the form
@@
for some nonnegative integer $r$, subsets $S_{1},S_{2},\ldots,S_{r}$ of $\left\{1,2,\ldots,d\right\}$ and nonzero coefficients $\alpha_{S_{j}},j=1,2,\ldots,d$. We say that $f^{*}$ satisfies the merged-staircase property (MSP) (Abbe et al., 2022) if the subsets $S_{1},S_{2},\ldots,S_{r}$ can be reordered such that for any $i\in\{1,2,\ldots,r\}$, we have
@@ 1.6 Robust Lower Bounds under “Soft” Bottlenecks
When $\mathcal{X}=\{\pm 1\}^{d}$, $\nu=\operatorname{Unif}\left(\{\pm 1\}^{d}\right)$, $f^{*}$ depends only on $s$ covariates, then the expected $L^{2}$ risk of the CART estimator $\hat{f}_{\operatorname{CART}}$ satisfies
$E_{\mathcal{D}_{n}}\left\{R\left(\hat{f}(-;\mathcal{D}_{n}),f^{*}\right)\right\}=\Omega(1),~{}\text{whenever}~{}n=\min\left\{O\left(1/\min_{f\notin\textsf{MSP}}R(f,f^{*})\right),\exp(O(d))\right\}.$
@@ 2.5 CART
where $\mathcal{I}(\mathcal{C};\mathcal{D}_{n})\coloneqq\frac{1}{N(\mathcal{C})}\sum_{i\in I(\mathcal{C})}\left(Y_{i}-\bar{Y}_{\mathcal{C}}\right)^{2}$ is called the impurity of the cell $\mathcal{C}$.
@@ 2.6 Random Forests
Slightly abusing notation, we use $\hat{f}_{\operatorname{CART}}(-;\mathcal{D}_{n},\theta)$ to denote the CART model fitted with a random seed $\theta$. A random forest model with $M$ trees is then
@@ 2.7 Greedy Tree Models
In order to isolate the aspect of CART that leads to its limitations with non-MSP functions, we make a more general definition. We say that $\mathcal{A}$ is a greedy tree algorithm if, similar to CART, it grows a tree greedily in a top-down fashion, but uses a possibly different criterion $\mathcal{O}(k;\mathcal{C},\mathcal{D}_{n})$ instead of impurity decrease (12) to determine its splits. We constrain $\mathcal{O}(k;\mathcal{C},\mathcal{D}_{n})$ to depend on $\mathcal{D}_{n}$ only via the values of the $k$-th covariate and the response for observations in $\mathcal{C}$, i.e.,
@@ 7.1 Characterizing High-Dimensional Consistency for CART
Next, one can prove that $f^{*}$ is SMSP if and only if it satisfies the SID condition (7) for some value of $\lambda>0$ (see Proposition F.2). Previous work, specifically Chi et al. (2022), who first introduced the SID condition, as well as Mazumder and Wang (2024), provide upper bounds for the estimation error of CART given the SID condition. These results and proofs are stated for continuous covariates, but can easily be adapted to a discrete covariate setting. We revisit this last statement in the next subsection. Assuming this for now, our discussion thus far yields the following characterization theorem.
@@
It is natural to ask whether our lower bounds can be extended to handle all $f^{*}$ that do not satisfy SMSP (or equivalently the SID condition). This turns out to be impossible—there exist functions that do not satisfy SMSP, but for which high-dimensional consistency holds. For instance, consider $f^{*}(\mathbf{x})=x_{1}+2x_{2}+2x_{1}x_{2}+x_{2}x_{3}$. Since $f^{*}|_{\left\{x_{1}=-1\right\}}=x_{2}x_{3}$, this does not satisfy SMSP, but on the other hand, CART would always first split on $x_{2}$ given enough samples, in which case this cancellation never occurs. Indeed, by reformulating CART as a sequential greedy approximation algorithm for convex optimization in a Hilbert space, Klusowski and Tian (2024) showed high-dimensional consistency for additive models and arbitrary covariates without the SID condition.
@@ 11 Discussion
Finally, while we have proved lower bounds for greedy axis-aligned regression trees, it is natural to ask whether lower bounds of a similar flavor hold for oblique regression trees as well as other more general classes of regression trees and ensembles under computational constraints. We hope our work will serve as a first step in clarifying the statistical-computational trade-offs inherent within the class of greedy algorithms fitting regression trees and ensembles.
@@ Acknowledgments
The authors would like to thank Yingying Fan, Samory Kpotufe, Wei-Yin Loh, Rahul Mazumder, and Denny Wu for insightful discussions. Tan gratefully acknowledges support from NUS Start-up Grant A-8000448-00-00 and MOE AcRF Tier 1 Grant A-8002498-00-00. Klusowski gratefully acknowledges financial support from the National Science Foundation (NSF) through CAREER DMS-2239448. Balasubramanian gratefully acknowledges support from NSF through grants DMS-2053918 and DMS-2413426.
@@ Appendix I Constant Probability Lower Bounds
The lower bounds portion of Abbe et al. (2023)’s conjecture is suggested by existing Statistical Query (SQ) and Correlational Statistical Query (CSQ) lower bounds. SQ and CSQ are generic models of constrained computation (Kearns, 1998; Feldman et al., 2021). In the SQ model, the analyst has access to a sequence of queries, each of which returns an approximation of the expectation $\mathbb{E}\left\{\psi(\textbf{X},Y)\right\}$ for some real-valued function $\psi$ that may change from query to query. The CSQ model further constrains the queries to be of the form $\mathbb{E}\left\{\phi(\textbf{X})Y\right\}$ for some real-valued function $\phi$. Under the SQ model, it is known that $\Omega\left(d^{k}\right))$ queries are required to learn $k$-parities (Barak et al., 2022), while Abbe et al. (2023) proved a CSQ lower bound of the same order. Now, Abbe et al. (2021) showed that SGD is no more powerful than the SQ model if the stochastic gradients have limited precision, but it is unclear to us whether this holds for the SGD regime in 2LNNOSGD. Assuming the reduction holds, then the SQ and CSQ lower bounds translate to a lower sample complexity bound of $\tilde{\Theta}\left(d^{k-1}\right)$ (Kou et al., 2024).
We also note an additional performance gap suggested by Abbe et al. (2023)’s conjecture. They in fact consider a generalization of sparse Boolean functions called multi-index functions. These are defined as $f^{*}(\mathbf{x})=g(\mathbf{M}\mathbf{x})$, where $\mathbf{M}$ is an unknown $d\times s$ matrix with orthogonal columns. In other words, $f^{*}$ depends only on an $s$-dimensional latent subspace. The fact that this subspace is not axis-aligned is irrelevant to SGD and neural networks, which are rotation-invariant. Glasgow (2024)’s upper bound for XOR was furthermore proved in this setting. [^8] Kou et al. (2024)’s result required axis alignment, which they exploited using sign SGD. On the other hand, having sparsity in the original covariate basis is crucial to the good performance of regression trees and ensembles in general. This is because every root-to-leaf path can split on at most $O(\log d)$ covariates, and hence cannot fit to any dense rotation of a latent subspace, no matter how small the latent dimension is (Tan et al., 2021). This drawback may be overcome using oblique (that is, non-axis-aligned) regression trees and ensembles (Cattaneo et al., 2024; O’Reilly, 2024), but these have much larger computational overhead. Finally, we note that for continuous covariates, the information exponent introduced by Arous et al. (2024) and its generalizations by Oko et al. (2024); Ren and Lee (2024) may be more effective in capturing learnability. Exploring similar measures of learnability for greedy recursive partitioning estimators presents an interesting direction for future research.